OpenStudy (anonymous):

how much energy doesw an office with 50 lightbulbs ( 40 W) each use every minute.. how fast would joe have to run to have the same amoiunt of kinetic energy?

OpenStudy (anonymous):

That WATT is named after the Scottish engineer James Watt (1736–1819). and is the expenditure of energy at the rate of 1J/s. You are required to determine the energy usage over the period of an hour. By multiplication you can determine what it is in one second so can proceed to determine what is is in one hour. You then need to "equate" this energy to to that which would be expended by a runner. What will the units of energy be? I'll leave it there and follow up hoping that this gets you moving on the problem.

OpenStudy (callisto):

W = Js^-1 40W = 40Js^-1 Total amount of energy consumed every minute= 50 x (40 x 60) = 120000J KE = (1/2)mv^2 120000 = (1/2)mv^2 240000 = mv^2 What is the mass of Joe?

OpenStudy (anonymous):

oh sorry the mass of joe is 15kg

OpenStudy (callisto):

240000 = mv^2 240000 = 15v^2 16000 = v^2 Can you find v? BTW Joe is too light :|

OpenStudy (anonymous):

yah it turns out to be 126

OpenStudy (callisto):

Perhaps this is the answer :|

OpenStudy (anonymous):

but the answer key says 16 ... i dont know why.. the first part of the question was corrrect

OpenStudy (anonymous):

it seems that for k 2000 was plugged in